INAF queue observing, 2013 May 9-10

Observer: F. Cusano, R. Speziali
Telescope Operator: S. Allanson
Support Astronomer: O. Kuhn
Instrument: LBC -MODS

Summary

In the afternoon snow, at the sunset clouds covering 50% of the sky.
At the beginning of the night during the LBC observations the sky was higly variable, with cirrus and clouds passing through.
We switch to MODS to collimate the secondary and try to observe Sommariva. Collimation not reached.
At 09:00 UT the sky is clear and we start observing with LBC.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Grazian_30A Goods_North Uspec, r 1h? We performed in total 1.5h of integration time, but the sky was variable (cirrus and clouds) during the observation
Brocato_32C Draco Uspec, I; B, V 17.5m x 6.6m completed
Photometric STD SA107 U, B, Uspec, g; V, r, i ,I 12m airmass 1.65
Piotto_34B M13 B, V 12.2m x 9.8m  
Calibration skyflat Uspec, V    

Details

03:10 The TO opens the dome

03:30 slew to Grazian30_A (seeing 1.1-1.2)

03:30 dofpia, /first, /X2

03:43 launched grazian_pointing.ob

03:48 image quality not good running dofpia, /X2

03:54 Launched grazian_GOODS_UR_37.ob stopped

03:55 dopfia, /X2

04:03 Launched grazian_GOODS_UR_37.ob (seeing 0.7 in the red, 0.8 in the blue, cirrus coming, from file 41995 and 042735 only bright stars in the frames)

04:50 dofpia, /X2

04:55 Launched again grazian_GOODS_UR_37.ob (seeing 0.7 both channels, some cirrus)

05:35 dofpia, /X2

05:38 Launched grazian_GOODS_UR_38.ob (seeing 1, clouds coming in)

06:25 Clouds covering almost all the sky

06:40 Switch to MODS to perform the collimation of the rigid secondary mounted this afternoon

09:05 Problem in the collimation of the secondary, switch to LBC

09:16 Slewing to Brocato32_C (Draco)

09:17 Launched Dra_pointing.ob

09:18 dopfia, /X2

09:29 lbcrangebal

09:33 Launched Dra.ob (filter Uspec,I; seeing 0.7 in both channels, airmass 1.1; integration time 17.5 minutes)

09:54 Launched Dra_2.ob (filter B,V; seeing 0.7; integration time 6.6 m)

10:14 Slew to standard field SA107

10:18 dofpia, /X2

10:20 Restarting LBCcontrol, in the blue channel there were no corrections for guiding and attive optics since 04:10 UT

10:31 dofpia on standard field SA107

10:36 Launched SA107_UBUsgVriI

10:53 Slew to Piotto_34B M13

10:53 dofpia, /X2

10:59 Launched M13_BV3.ob (seeing 1.0)

11:08 Launched M13_BV4.ob

11:17 Launched M13_BV5.ob (twilight)

11:33 Starting skyflat test

11:36 error in the red filter wheel, reboot LBCcontrol

11:43 Starting skyflat test

11:48 skyflat_SV x5, too bright we try x3 and x1

11:59 skyflat_BI x0.3 too brigth

12:04 closing dome

-- FeliceCusano - 10 May 2013
Topic revision: r5 - 12 May 2013, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback